OpenStudy (anonymous):

If the distance between two points on the coordinate axes is the square root of 65 , which of the following could be the coordinates of the two points?

OpenStudy (p0sitr0n):

\[distance=\sqrt{ (\Delta x)^{2} + (\Delta y)^{2} } \]

OpenStudy (p0sitr0n):

where delta x = x2 -x1 delta y =y2 -y1